OpenStudy (anonymous):

If f(x) = x/2 + 8, what is f(x) when x = 10?

OpenStudy (anonymous):

in need help

OpenStudy (anonymous):

in need help

Nnesha (nnesha):

just substitute x for its given value

OpenStudy (anonymous):

ok but theses are my answer choices4 9 13 36

Nnesha (nnesha):

okay. given x=10 so substitute x for 10

OpenStudy (keaganizaac):

(10)/2 = 5 + 8 = 13

OpenStudy (anonymous):

so whats the answer

Nnesha (nnesha):

how would it look like after substitutin x for 10 ?and then reduce the fracation what would you get ??

OpenStudy (anonymous):

If f(x) = 6x – 4, what is f(x) when x = 8?

OpenStudy (keaganizaac):

6(8) = 48 - 4 = 44

OpenStudy (anonymous):

If f(x) = 5x + 40, what is f(x) when x = –5?

OpenStudy (keaganizaac):

5(-5) + 40. -25 + 40, 15

OpenStudy (anonymous):

What is the inverse of the function f(x) = x 1/9+ 2?

OpenStudy (keaganizaac):

hold on

OpenStudy (keaganizaac):

did you mean 1/9x + 2?

OpenStudy (anonymous):

yeah whats the inverse of the function

OpenStudy (keaganizaac):

Ok, sorry i didnt realize how many answers i have given and Im getting notifications from moderators, its a simple question, if the line has a slope of 1/9, how would you make it the opposite?